A piece of machinery costs $367,000 and has an anticipated $35,000 resale value at the end of its five-year useful life. When preparing a depreciation schedule for the piece of machinery by the-sum-of-the-years method, what is the second year book value? (Answer with an integer; Round to the nearest integer)
Answer #1

Sum of years = n(n+1)/2 = 5(5+1)/2 = 15

Depreciation for year 1 = (367,000-35,000) * 5/15 = 110,667

Second Year book value = Cost of asset - Depreciation for year 1

= 367,000 - 110,667

= 256,333